### v2

# Last Update: September, 5, 2024

Last Update: September, 5, 2024
Differentially Private Kernel Density Estimation
Erzhi Liu111 erzhiliu@u.northwestern.edu. Center for Foundation Models and Generative AI & Department of Computer Science, Northwestern University, Evanston, IL 60208, USA  Jerry Yao-Chieh Hu222jhu@ensemblecore.ai; jhu@u.northwestern.edu. Ensemble AI, San Francisco, CA 94133, USA; Center for Foundation Models and Generative AI & Department of Computer Science, Northwestern University, Evanston, IL 60208, USA Alex Reneau333alex@ensemblecore.ai. Ensemble AI, San Francisco, CA 94133, USA  Zhao Song444magic.linuxkde@gmail.com. Simons Institute for the Theory of Computing, UC Berkeley, Berkeley, CA 94720, USA  Han Liu555hanliu@northwestern.edu. Center for Foundation Models and Generative AI & Department of Computer Science & Department of Statistics and Data Science, Northwestern University, Evanston, IL 60208, USA
**footnotetext: These authors contributed equally to this work. This work was done during JH’s internship at Ensemble AI.
We introduce a refined differentially private (DP) data structure for kernel density estimation (KDE) with $\ell_{1}$, $\ell_{2}$ and $\ell_{p}^{p}$ kernels. This new DP data structure offers not only improved privacy-utility tradeoff but also better query efficiency over prior results. Specifically, we study the mathematical problem: given a similarity function $f$ (or DP KDE) and a private dataset $X\subset\mathbb{R}^{d}$, our goal is to preprocess $X$ so that for any query $y\in\mathbb{R}^{d}$, we approximate $\sum_{x\in X}f(x,y)$ in a differentially private fashion. The best previous algorithm for $f(x,y)=\|x-y\|_{1}$ is the node-contaminated balanced binary tree by [Backurs, Lin, Mahabadi, Silwal, and Tarnawski, ICLR 2024]. Their algorithm requires $O(nd)$ space and time for preprocessing with $n=|X|$. For any query point, the query time is $d\log n$, with an error guarantee of $(1+\alpha)$-approximation and $\epsilon^{-1}\alpha^{-0.5}d^{1.5}R\log^{1.5}n$.
In this paper, we use the same space and pre-processing time, improve the best previous result [Backurs, Lin, Mahabadi, Silwal, and Tarnawski, ICLR 2024] in three aspects
  • We reduce query time by $\alpha^{-1}\log n$ factor
  • We improve the approximation ratio from $\alpha$ to $1$
  • We reduce the error dependence by a factor of $\alpha^{-0.5}$
From a technical perspective, our method of constructing the search tree differs from previous work [Backurs, Lin, Mahabadi, Silwal, and Tarnawski, ICLR 2024]. In prior work, for each query, the answer is split into $\alpha^{-1}\log n$ numbers, each derived from the summation of $\log n$ values in interval tree countings. In contrast, we construct the tree differently, splitting the answer into $\log n$ numbers, where each is a smart combination of two distance values, two counting values, and $y$ itself. We believe our tree structure may be of independent interest.
1 Introduction
We propose a refined differentially private (DP) data structure for DP kernel density estimation, offering improved privacy-utility tradeoff over prior results without compromising efficiency. Let $X\subset\mathbb{R}^{d}$ be a private dataset, and let $f(x,y):\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a similarity function [^1] In this work, we use “kernel query” and “distance query” interchangeably., such as a kernel or distance function, between a user query $y\in\mathbb{R}^{d}$ and a private data point $x\in X$.
Problem 1 (DP Kernel Density Estimation Query).
The DP Kernel Density Estimation (KDE) query problem aims for an algorithm outputting a private data structure $D_{X}:\mathbb{R}^{d}\rightarrow\mathbb{R}$, that approximates the map $y\mapsto\sum_{x\in X}f(x,y)$ in a DP fashion [^2] For a given dataset $X$ and a given query $y$, an KDE query computes an approximation to $\sum_{x\in X}f(x,y)$. . Especially, we require $D_{X}$ to remain private with respect to $X$, regardless of the number of queries.
Problem 1 is well-studied [HRW13, HR14, WJF+16, AR17, ACSS20, CS21, AA22, QRS+22, AS23, GSY23, WNM23, AS24, HLSL24, LLSS24, BLM+24] due to its broad applicability and its importance in productizing large foundation models [LGK+24, XLB+24]. What makes this problem interesting is its generic abstraction of many practical DP challenges in modern machine learning. These include generating synthetic data similar to a private dataset [LJW+20, LTLH22, YNB+22, YLJ+22], and selecting a similar public dataset for pre-training ML models [HZS+23, YGK+24, YIL+23]. Such problem becomes more prevalent in this era of large foundation models [LGK+24, XLB+24]. Essentially, these problems involve computing the similarity between a private dataset (i.e., $X$) and a processed data point (i.e., query $y$), thus falling under Problem 1.
In this note, we focus on the $\ell_{1}$ kernel [^3] We remark that $\ell_{1}$ query distance is fundamental. Namely, it is easy to generalize the $\ell_{1}$ results to a wide range of kernels and distance functions following the provably dimensional reduction recipe of [BLM+24]. See Appendix B and C for improved DP KDE results with $\ell_{2}$ and $\ell_{p}^{p}$ kernels. (i.e., $f(x,y)=\|x-y\|_{1}$). By far, the algorithm with the best privacy-utility tradeoff and query efficiency for Problem 1 is by Backurs, Lin, Mahabadi, Silwal, and Tarnawski [BLM+24]. They propose a DP data structure via a node-contaminated balanced binary tree for $\ell_{1}$ kernel. To be concrete, we begin by defining the similarity error between two data structures and then present their results.
Definition 1.1 (Similarity Error between Two Data Structures).
For a fixed query, let $A$ represent the value output by our private data structure, and let $A^{\prime}$ represent the true value. We say that $A$ has an error of $(M,Z)$ for $M\geq 1$ and $Z\geq 0$, if $\mathbb{E}[|A-A^{\prime}|]\leq(M-1)A^{\prime}+Z$. This implies a relative error of $M-1$ and an additive error of $Z$. The expectation is taken over the randomness used by the data structure.
Let $n:=|X|$ be the size of the dataset, $X_{i,j}$ be the $j$-th dimension of the $i$-th data point, $R:=\max_{i,j}(X_{i,j})$ be the max value of each data entries from each dimension, and $\alpha\in[0,1]$ be a parameter of the data structure selected before calculation. For $\ell_{1}$ kernel, [BLM+24] gives the error bound $(1+\alpha,\alpha^{-0.5}\epsilon^{-1}Rd^{1.5}\log^{1.5}n)$ through their DP data structure.
Compared to the results of [BLM+24], we provide a refined DP data structure and improves both privacy-utility tradeoff and query efficiency. Specifically, we not only improve the $\ell_{1}$ error bound to $(1,\epsilon^{-1}Rd^{1.5}\log^{1.5}n)$, but also the query time from $O(\alpha^{-1}d\log^{2}n)$ to $O(d\log n)$. This makes our new DP data structure the latest best algorithm for Problem 1:
Theorem 1.1 (Informal Version of Theorem 3.2).
Given a dataset $X\subset\mathbb{R}^{d}$ with $|X|=n$. There is an algorithm that uses $O(nd)$ space to build a data-structure which supports the following operations:
  • Init$(X,\epsilon)$. It takes dataset $X$ and privacy parameter $\epsilon$ as input and spends $O(nd)$ time to build a data-structure.
  • Query$(y\in\mathbb{R}^{d})$. It takes $y$ as input and spends $O(d\log n)$ time to output a scalar $A$ such that $\mathbb{E}[|A-A^{\prime}|]\leq O(\epsilon^{-1}Rd^{1.5}\log^{1.5}n)$.
Notably, our improvements stem from a key observation that — in the balanced binary tree data structure proposed by [BLM+24], each query answer is split into $\alpha^{-1}\log n$ values, derived from the summation of $\log n$ values in interval tree counts. We deem that this splitting is not essential and can be achieved through simple and efficient preprocessing steps.
To this end, we introduce a novel data representation in the tree nodes, where each node stores the sum of distances from one point to multiple points. This allows us to split the answer into only $\log n$ values, each being a smart combination of two distance values, two count values, and $y$ itself. This additional information enhances each query calculation, leading to an improved privacy-utility tradeoff (Theorem 3.2 and Lemma 3.4) and faster query time (Lemma 3.3).
References Approx. Ratio Error Query Time Init time Space
[BLM+24] $1+\alpha$ $\alpha^{-0.5}\epsilon^{-1}Rd^{1.5}\log^{1.5}n$ $\alpha^{-1}d\log^{2}n$ $nd$ $nd$
Theorem 1.1 $1$ $\epsilon^{-1}Rd^{1.5}\log^{1.5}n$ $d\log n$ $nd$ $nd$
Table 1: Comparison of $\norm{x-y}_{1}$ Results with Best Known Algorithm by [BLM+24]. We ignore the big-O notation in the table, for simplicity of presentation. We use $n$ to denote the number of points in dataset. We use $d$ to denote the dimension for each point in the dataset. We assume all the points are bounded by $R$. We use the $\epsilon$ to denote $\epsilon$-DP.
Lastly, we remark that our $\ell_{1}$ results are transferable to other kernels via the provably dimensional reduction recipe of [BLM+24]. To showcase this, we generalize our results to $\ell_{2}$ and $\ell_{p}^{p}$ kernels following [BLM+24]. Notably, these results also improve upon the best previous results reported in [BLM+24] for $\ell_{2}$ and $\ell_{p}^{p}$ with $p\in[1,n]$. For details, see Appendix B and C, particularly the comparison Tables 2 and 3.
2 Preliminaries
For a positive integer $n$, we use $[n]$ to denote $\{1,2,\cdots,n\}$. For a vector $x$, we use $\|x\|_{1}:=\sum_{i=1}^{n}|x_{i}|$ to denote its $\ell_{1}$-norm. We use $\mathbb{E}[\cdot]$ to denote the expectation. We use $\mathop{\mathrm{Var}}[\cdot]$ to denote the variance. We use $\Pr[\cdot]$ to denote the probability. We use $\mathsf{Laplace}(\lambda)$ random variable with parameter $\lambda$. It is known that $\mathbb{E}[\mathsf{Laplace}(\lambda)]=0$ and $\mathop{\mathrm{Var}}[\mathsf{Laplace}(\lambda)]=2\lambda^{2}$.
2.1 Basic Facts
Fact 2.1.
Let $X$ denote a random variable with $\mathbb{E}[X]=0$. Then it holds
$\mathbb{E}[|X|]\leq(\mathop{\mathrm{Var}}[X])^{1/2}.$
Proof.
It is easy to see that $(\mathbb{E}[|X|])^{2}\leq\mathbb{E}[|X|^{2}]=\mathbb{E}[X^{2}]$. Next, we show
$(\mathbb{E}[|X|])^{2}\leq\leavevmode\nobreak\ \mathbb{E}[X^{2}]=\leavevmode\nobreak\ \mathbb{E}[X^{2}]-(\mathbb{E}[X])^{2}\leq\leavevmode\nobreak\ \mathop{\mathrm{Var}}[X].$
Thus, we complete the proof. ∎
2.2 Differential Privacy
For completeness, we state the definitions of differential privacy [DMNS06, DRV10].
Definition 2.1 (Pure/Approximate Differential Privacy).
A randomized algorithm $M:\mathcal{X}^{n}\rightarrow\mathcal{Y}$ satisfies ($\epsilon$, $\delta$)-differential privacy if, for all $x,x^{\prime}\in\mathcal{X}^{n}$ differing on a single element and for all events $E\subset\mathcal{Y}$, we have $\Pr[M(x)\in E]\leq e^{\epsilon}\cdot\Pr[M(x^{\prime})\in E]+\delta$
For special case of ($\epsilon$, 0)-differential privacy, we use pure or pointwise $\epsilon$-differential privacy to denote it. For the other case of $\delta$ > 0, we denote it as approximate differential privacy.
Next, we introduce a common tool in DP proofs. For multiple independent DP functions, the next lemma provides an estimation on the privacy of their composition:
Lemma 2.1 (Advanced Composition Starting from Pure DP [DRV10]).
Let $M_{1},...,M_{k}:X^{n}\rightarrow Y$ be randomized algorithms, each of which is $(\epsilon,\delta)$-DP. Define $M:X^{n}\rightarrow Y^{k}$ by $M(x)=(M_{1}(x),...,M_{k}(x))$ where each algorithm is run independently. Then $M$ is $(\epsilon^{\prime},\delta)$-DP for any $\epsilon,\delta>0$ and
$\epsilon^{\prime}=k\epsilon^{2}/2+\epsilon\sqrt{2k\log(1/\delta)}$
For $\delta=0$, $M$ is $k\epsilon$-DP.
2.3 [BLM+24]’s DP Data Structure: Node-Contaminated Balanced Tree
[BLM+24] propose a data structure for general high-dimensional (e.g., $d$-dimensional) $\ell_{1}$ kernel via $1$-dimensional decomposition: $\sum_{x\in X}\|x-y\|_{1}=\sum_{i=1}^{d}\sum_{x\in X}|x_{i}-y_{i}|$. Namely, they create a DP data structure for a $d$-dimensional dataset $X$ by considering $d$ copies $1$-dimensional DP data structures (i.e., $\sum_{x\in X}|x_{i}-y_{i}|$ for $i\in[d]$). For each of these $1$-dimensional DP data structures, they employ a node-contaminated balanced tree algorithm as follows.
Let $n=|X|$ be the size of the dataset $X$. Considering all input values are integer multiples of $R/n$ in $[0,R]$, [^4] For general continuous data, they achieve this by rounding all data points to some integer multiples of $R/n$.
  1. 1.
    They use a classic balance binary tree of $L$ layers in one dimension: a binary tree where each non-leaf node has exactly two children, and all leaf nodes are at the same depth, $L$. The tree has $12^{L}-$ 1 total nodes, with $2^{L-1}$ leaf nodes and $L$ levels from the root (level $1$) to the leaves (level $L$). Each level $k$ contains $2^{k-1}$ nodes.
  2. 2.
    They assign the leaves to correspond to these multiples (of $R/n$ in $[0,R]$) and store the number of dataset points at each position, while internal nodes store the sum of their children.
  3. 3.
    It’s well-known that any interval query can be answered by summing values from $O(\log n)$ tree nodes. This provide an efficient (noise-less) query operation.
  4. 4.
    To ensure DP, they propose a noisy version of the tree by injecting noise onto each node, i.e., a node-contaminated balanced tree. By above, changing any data point affects only $O(\log n)$ counts in the tree, each by at most one (from leaf to root), which bounds the sensitivity of the data structure.
To highlight the significance of this work — removing the $\alpha$ dependence in the result and eliminating a $\log n$ factor in the query (see Table 1) — we make the following remarks on [BLM+24].
  • Query Time. For each query $y$, in order to report the answer, they need to take summation of $O(\alpha^{-1}\log n)$ numbers where each number from each region $(y+(1+\alpha)^{i},y+(1+\alpha)^{i+1}]$ (see line 7 in Algorithm 3 in [BLM+24]). In particular, for each region, they need to run a interval query which consists of $\log n$ numbers (each number is from a tree node, see Algorithm 2 in [BLM+24]). That’s why the their algorithm has $\alpha^{-1}\log^{2}n$ dependence.
  • Error Bound and Accuracy. Due to the region $(y+(1+\alpha)^{i},y+(1+\alpha)^{i+1}]$ can only approximate the number within $(1+\alpha)$ approximation. That is why the final error is proportional to $\alpha^{-1}$ and accuracy is losing a $(1+\alpha)$ relative error. Since the values are in range $[0,R]$, thus the error bound will be linearly depend on $R$ according to how they construct the tree and alayze the error.
  • Sensitivity and DP Guarantee. Since the tree is analyzing the counting problem, thus the sensitivity is $O(\log n)$ because change one point, will affect $O(\log n)$-levels stored counts.
2.4 High-Level Overview of Our DP Data Structure
We improve [BLM+24] by providing a refined balanced tree for DP $1$-dimensional $\ell_{1}$ KDE. Specifically, we introduce a new data representation on each node of the balanced binary tree as follows.
Let $n=|X|$ be the size of the dataset $X$ and $\{x_{k}\}\in[0,R)$ for $k\in[n]$.
  1. 1.
    We use a balanced binary tree where in each node stores both distance $s$ and the counts $c$. Here, $s$ denotes the pre-summation of certain distances, and $c$ denotes the pre-counting.
  2. 2.
    We make a key observation (Lemma 3.1) that these values ($s$ and $c$) are not depending on query $y$, while they are critical components of the final answer (the $\ell_{1}$ kernel $\sum_{x\in X}\|x-y\|_{1}$). This motivate us to construct the answer by using $L=\log n$ layers result, where each layer result is a smart combination of $y,s_{1},s_{2},c_{1},c_{2}$ (i.e., $(s_{1}-s_{2}+y\cdot c_{1}-y\cdot c_{2})$ ).
  3. 3.
    To ensure DP, we need a noise version of the tree, due to $s$ and $c$ have different sensitivities, thus we need to add different levels of noise for $s$ and $c$.
A few remarks are in order:
  • Query Time. For each query $y$, to report the answer, we first extract four $L=O(\log n)$ numbers along with $y$ and apply a special function: $(s_{1}-s_{2}+y\cdot(c_{1}-c_{2}))$ (see our key observation in Lemma 3.1).
  • Error Bound and Accuracy. Since our algorithm does not use the $(1+\alpha)^{i}$ region concept, we avoid the $\alpha^{-1}$ error. As a result, our error is purely additive, with no relative error.
  • Sensitivity and DP Guarantee. Since $s$ approximates the summation of distances, the sensitivity is $O(LR)$, where each node in one of the $L$ layers contributes a factor of $R$. On the other hand, since $c$ counts the points, its sensitivity is $O(L)$, with each node in the $L$ layers contributing an $O(1)$ factor. Previous work [BLM+24] only use $y$ to determine which region to count points, but our algorithm uses $y$ to construct the final answer. Since $y\leq R$, this explains the $R$ gap between the noise levels in $s$ and $c$.
2.5 Related Works
Privacy, Security and Safety in Large Foundation Models.
The motivating problems for this work come from the privacy concerns of large foundation models [LGK+24, XLB+24]. In modern machine learning, Foundation models [BHA+21], including pretrained transformer and diffusion models, gain popularity in many AI applications due to their ability to generalize across diverse tasks with minimal fine-tuning. Pretrained transformers, such as BERT [DCLT18], GPT [BMR+20], and Llama [TMS+23, TLI+23], leverage vast amounts of data to learn general-purpose, context-aware representations. Diffusion models [PX23, HJA20, SE19], on the other hand, excel in generative tasks, particularly in producing high-quality images and data distributions through iterative refinement [NDR+21, RDN+22, LZL+24, ZCY+24, ZXC+24, WZY+24, WWS+24]. Importantly, what makes them fundamental and versatile is their flexibility for diverse downstream tasks via fine-tuning methods [ZZZ+24, DQY+22, LARC21, LJF+21, HysW+22], hence “foundation.” Together, these models signify a shift towards more powerful AI systems that serve as the basis for a wide range of applications across different domains.
However, despite their empirical [FC20, YLW23, WIL+23, NPF+24, ZWH+24, ZJL+23, JZLD21, TTE+23, SAT+23, MBA+23, HCL+24, LZL+24, EKB+24, MGA+24, PX23, GZCY23, ZNVA23, BLCZ22] and theoretical successes [BCW+23, WCWH23, KS24, ZL24, LWL+24, LWLC23, LIPO23, TLTO23, WHHL24, WHL+24, AS24, AS23, HSK+24, HYW+23, RSL+20, YBR+20, LLSS24, SWXL24, XSL24, LSSZ24b, LSSZ24a, LSSS24, LLS+24b, LLS+24a, LSSY24, WMS+24, CMFW24, FDG+24, FYWC24, WCL+24, DCWY24, GYY+24, HWL+24, LWCC24, ZLC23, WWY24, CCL+23a, CHZW23, OAS23, CCL+23b], there exists a gap from productizing many these advancements due to privacy, safety, and security concerns [QHZ+24, SHW+24, LHLW24, WMR+21]. These models, due to their vast scale and extensive training on large datasets, risk exposing sensitive information and amplifying biases present in the data. Privacy issues arise when models inadvertently memorize and reproduce personal data from training sets [CHN+23, CTW+21]. Safety concerns include the potential for generating harmful or misleading content, especially in applications where accuracy is paramount [WMR+21, Ope23, TAB+23, Ant23, YLH+24, LYZ+24, SCB+24, DLL+23, LDX+23, LXCX23, SPT+23, YLX23]. Security vulnerabilities also exist, as these models may be susceptible to adversarial attacks that manipulate outputs or extract proprietary information [ZYZ+24, JXN+24, XQZ+24]. Addressing these challenges is essential to ensure the responsible and secure use of foundation models across various applications. In this work, we focus on the privacy of foundation models at a fundamental level by study the formalized KDE query Problem 1.
Differential Privacy (DP).
Differential Privacy (DP) is a standard tool for understanding and mitigating privacy concerns in machine learning. Proposed by [DMNS06], DP offers a robust approach to protecting sensitive information. Given the fact that non-private ML models expose sensitive user data [FJR15, CLE+19, CYZF20, CTW+21, CCN+22, CIJ+23, HVY+22, TSJ+22, CHN+23], researchers propose various methodologies. One approach involves generating synthetic datasets that closely resemble the original private data and training models on these synthetic datasets [LJW+20, LTLH22, YNB+22, YLJ+22, YIL+23, LGK+24]. Another strategy is to use similar public examples for pre-training models [HZS+23, YGK+24]. Additionally, the DP-SGD method has benefited from incorporating public data similar to private datasets to enhance downstream model performance [YZCL21, DBH+22, YNB+22, YLJ+22, LTLH22, HZS+23, YGK+24, LGK+24]. As highlighted by [BLM+24], the common denominator of all these works is the need to compute similarities to a private dataset. In this work, we study this problem at a fundamental level by formalizing it as the PD-KDE query problem (Problem 1).
Kernel Density Estimation (KDE).
Kernel Density Estimation (KDE) is a key technique in statistics and machine learning. This technique converts a collection of data points into a smoothed probability distribution [SS02, STC04, HSS07]. It is prevalent in many private applications such as crowdsourcing and location sharing [HWM+19, CCF21]. Although research on non-private KDE already produce efficient methods [BCIS18, BIW19, ACSS20, CKNS20, LSXZ22, QRS+22, BIK+22, HSW+22, DJS+22], adapting these techniques to DP remains complex [GRU12, BLR13, AR17, WNM23, BLM+24]. The best algorithm to date is by [BLM+24]. In this work, we introduce a new data structure for DP-KDE that improves upon [BLM+24] in both privacy-utility trade-off and efficiency.
3 A Refined Differentially Private Data Structure
We propose a new data structure with $\epsilon$-differential privacy that achieves $O(\epsilon^{-1}R\log^{2}n)$ additive error. Our method improves the balanced binary tree structure proposed by Backur, Lin, Mahabadi, Silawal, and Tarnawsk [BLM+24] in both DP guarantee and efficiency. Our data structure adapts a new data representation on each node that simplifies each query without losing privacy guarantees.
3.1 Key Observation and New Data Structure for One Dimensional $\ell_{1}$ KDE Query
For simplicity, we consider the 1-dimensional KDE query problem with $\ell_{1}$ kernel distance. Note that any high-dimensional KDE query problem can be reduced to this case via the 1-dimensional decomposition discussed in Section 2.3 or in [BLM+24].
When doing queries, we observe that
Lemma 3.1.
For a collection of values $\{x_{1},x_{2},\cdots,x_{n}\}\subset\mathbb{R}$ and a value $y$, we define two sets
$S_{+}:=$$\leavevmode\nobreak\ \{k\in[n]\leavevmode\nobreak\ :\leavevmode\nobreak\ x_{k}>y\}$
$S_{-}:=$$\leavevmode\nobreak\ \{k\in[n]\leavevmode\nobreak\ :\leavevmode\nobreak\ x_{k}<y\}.$
It holds
$\sum_{k=1}^{n}|x_{k}-y|=\underbrace{(\sum_{k\in S_{+}}x_{k})}_{s_{\rm right}}-\underbrace{(\sum_{k\in S_{-}}x_{k})}_{s_{\rm left}}+y\cdot\underbrace{|S_{-}|}_{c_{\rm left}}-y\cdot\underbrace{|S_{+}|}_{c_{\rm right}}.$
Proof.
$\sum_{k=1}^{n}|x_{k}-y|=$ $\leavevmode\nobreak\ \sum_{k\in S_{+}}(x_{k}-y)+\sum_{k\in S_{-}}(y-x_{k})$
$=$ $\leavevmode\nobreak\ (\sum_{k\in S_{+}}x_{k})-(\sum_{k\in S_{-}}x_{k})+y\cdot(\sum_{k\in S_{-}}1)-y\cdot(\sum_{k\in S_{+}}1)$
$=$ $\leavevmode\nobreak\ (\sum_{k\in S_{+}}x_{k})-(\sum_{k\in S_{-}}x_{k})+y\cdot|S_{-}|-y\cdot|S_{+}|.$
This completes the proof. ∎
Lemma 3.1 provides a neat decomposition of the $\ell_{1}$ KDE query into four components. This motivates us to design a new PD data structure that pre-computes these terms and storing them in a balanced tree for possible algorithmic speedup. Surprisingly, this preprocessing not only accelerates query time but also improves the privacy-utility tradeoff. The following discussion illustrates this.
To calculate each part efficiently, we define our data representation on a balanced binary tree as follows:
  • Dataset: Given a dataset $X\coloneqq\{x_{k}\}_{k=1}^{N}$ containing $n$ values in the range $[0,R)$, we build a balanced binary tree using $X$.
  • Tree Structure:
    • Let $L$ denote the total number of layers in the tree.
    • At the $l$-th layer, there are exactly $2^{l}$ nodes.
  • Node Representation:
    • Each node represents a consecutive interval of $X$.
    • The interval for the $j$-th node at the $l$-th layer is defined as $I_{l,j}:=[(j-1)\cdot\frac{R}{2^{l-1}},j\cdot\frac{R}{2^{l-1}})$.
    • The left and right children of $I_{l,j}$ are $I_{l+1,2j-1}$ and $I_{l+1,2j}$, respectively.
Each node $I_{l,j}$ stores two values: $c_{l,j}$ and $s_{l,j}$. Here:
  • $c_{l,j}:=|\{x_{k}:x_{k}\in I_{l,j}\}|$ represents the count of data points in the interval.
  • $s_{l,j}:=\sum_{x_{k}\in I_{l,j}}x_{k}$ represents the sum of the data points’ values in the interval.
After calculating each $c_{l,j}$ and $s_{l,j}$, we add independent noise drawn from $\mathsf{Laplace}(L/\epsilon)$ to each $c$ value and from $\mathsf{Laplace}(LR/\epsilon)$ to each $s$ value. The Init algorithm (Algorithm 1) outlines the construction of our data structure.
In the Query algorithm (Algorithm 2), we sum the $s$ and $c$ values at the relevant nodes. Here:
  • $s_{\rm left}$ represents $\sum_{k\in S_{-}}x_{k}$
  • $c_{\rm left}$ represents $|S_{-}|$
  • $s_{\rm right}$ represents $\sum_{k\in S_{+}}x_{k}$
  • $c_{\rm right}$ represents $|S_{+}|$
For each query, we first locate the leaf that $y$ belongs to, then traverse the tree from the bottom up, summing the values of the left and right nodes.
Next, we prove the time complexity, differential privacy property and error of results of out algorithm respectively.
Algorithm 1 Building Binary Tree
1:data structure FasterTree $\triangleright$ Theorem 3.1
2:members
3:     $T:=\{c,s\}$triangleright $c_{i,j}$ and $s_{i,j}$ represents the count and sum of node defined above
4:end members
5:procedure Init($X,n,\epsilon,L$) $\triangleright$ Lemma 3.2, Lemma 3.4
6:     for each $x_{k}$ in $X$ do
7:         find the leaf node $j$ of layer $L$ that $x_{k}$ is in the interval $I_{L,j}=[(j-1)\cdot\frac{R}{2^{L-1}},j\cdot\frac{R}{2^{L-1}})$
8:         $c_{L,j}\leftarrow c_{L,j}+1$
9:         $s_{L,j}\leftarrow s_{L,j}+x_{k}$
10:     end for
11:     for each layer $l$ from $L-1$ to $1$ do
12:         for each node $j$ in layer $l$ do
13:              $c_{l,j}\leftarrow c_{l+1,2j-1}+c_{l+1,2j}$
14:              $s_{l,j}\leftarrow s_{l+1,2j-1}+s_{l+1,2j}$
15:         end for
16:     end for
17:     Add independent noises drawn from $\mathsf{Laplace}({LR}/{\epsilon})$ to each $s_{l,j}$
18:     Add independent noises drawn from $\mathsf{Laplace}({L}/{\epsilon})$ to each $c_{l,j}$
19:     $T\leftarrow\{c,s\}$
20:end procedure
21:end data structure
Algorithm 2 One Dimensional Distance Query
1:data structure FasterTree $\triangleright$ Theorem 3.1
2:procedure Query($y\in\mathbb{R}$) $\triangleright$ Lemma 3.3, Lemma 3.5
3:     $s_{\rm left},c_{\rm left},s_{\rm right},c_{\rm right}\leftarrow 0$triangleright As previously defined
4:     for each layer $l$ from $2$ to $L$ do
5:         find the node $j$ of layer $l$ that $y$ is in the interval $I_{l,j}=[(j-1)\cdot\frac{R}{2^{l-1}},j\cdot\frac{R}{2^{l-1}})$
6:         if $j\mod 2=0$ then $\triangleright$ if node $j$ is the right child of its parent
7:              $c_{\rm left}\leftarrow c_{\rm left}+c_{l,j-1}$
8:              $s_{\rm left}\leftarrow s_{\rm left}+s_{l,j-1}$
9:         else $\triangleright$ if node $j$ is the left child of its parent
10:              $c_{\rm right}\leftarrow c_{\rm right}+c_{l,j+1}$
11:              $s_{\rm right}\leftarrow s_{\rm right}+s_{l,j+1}$
12:         end if
13:     end for
14:     return $s_{\rm left}-s_{\rm right}+y\cdot c_{\rm right}-y\cdot c_{\rm left}$triangleright Lemma 3.1
15:end procedure
16:end data structure
3.2 Time Complexity
Here we compute the init and query time of Algorithm 1 and 2, respectively.
Lemma 3.2 (Init Time).
If the total layer $L=\log(n)$, the running time of Init (Algorithm 1) is $O(n)$.
Proof.
By definition, on $l$-th layer, there are $2^{l}$ nodes. Therefore, the total number of nodes is $\sum_{l=1}^{L}2^{l}=2^{L+1}-1$. When $L=\log(n)$, the total number of nodes on the tree is $O(n)$. In Init (Algorithm 1), we iterate over $n$ data points, then iterate all nodes. Therefore the total time complexity of pre-processing is $O(n)$. ∎
Lemma 3.3 (Query Time).
If the total layer $L=\log(n)$, the time Query function (Algorithm 2) $O(\log(n))$ time.
Proof.
For each single query, we iterate the node that $y$ belongs to on each layer. The total layer number is $\log(n)$, so the time complexity of each query is $O(\log(n))$. ∎
3.3 Privacy Guarantees
Here we provide our DP guarantee of our proposed DP data structure (i.e., Algorithm 1).
Lemma 3.4 (Differential Privacy).
The data structure FasterTree returned by FasterTree.Init (Algorithm 1) is $\epsilon$-DP.
Proof.
For binary tree of $L$ layers, there are $12^{L+1}-$ 1 nodes. On each node, there are $2$ values $s$ and $c$. We consider $c$ and $s$ separately.
Let the functions $F_{c}(X)$ and $F_{s}(X):[0,R)^{n}\rightarrow\mathbb{R}^{2^{L+1}-1}$ represent the mappings from dataset $X$ to $c$ and $s$, respectively. Next, we prove that both $F_{c}(X)$ and $F_{s}(X)$ are $\epsilon/2$-DP.
When changing each data point, only the nodes containing the point change their values. On each layer, there is at most one such node. Therefore, only $O(L)$ values change.
Sensitivity for Storing Counts. For $F_{c}$, each $c$ value changes by at most $1$, so the sensitivity is $L$. Adding coordinate-wise Laplace noise with magnitude $\eta=O(2L/\epsilon)$ suffices to ensure $\epsilon/2$-DP using the standard Laplace mechanism.
Sensitivity for Storing Distances. For $F_{s}$, changing one data entry affects at most $O(L)$ values, and each value can be affected by at most $R$. Therefore, the sensitivity is $RL$. Adding coordinate-wise Laplace noise with magnitude $\eta=O(2RL/\epsilon)$ ensures $\epsilon/2$-DP.
By Lemma 2.1, the differential privacy parameter $\epsilon$ of the tree $T:=(F_{c}(X),F_{s}(X))$ equals $2\cdot\epsilon/2=\epsilon$. This completes the proof. ∎
3.4 Error Guarantee
Here we provide the (data-structure) similarity error of our proposed DP data structure.
Lemma 3.5 (Error Guarantee).
Let $X=\{x_{i}\}_{i\in[n]}$ be a dataset of $n$ one dimensional values $x_{i}\in[0,R)$. Let $A^{\prime}=\sum_{i=1}^{n}|x_{i}-y|$ be the true distance query value. Let $A$ be the output of Query (Algorithm 2) with $L=\log(n)$. Then we have $\mathbb{E}[|A-A^{\prime}|]\leq O(\log^{1.5}(n)R/\epsilon)$.
Proof.
We analyze the additive error of our algorithm. We notice that the difference between true distance and our output can be divided into two parts.
Part I.
The first part is the data in the leaf node which contains query $y$. In Query(Algorithm 2), We ignore these data points. Let the leaf node $i$ of layer $L$ be the node that query point $y$ belongs to. The error is
$\sum_{x_{k}\in[(j-1)\cdot R/2^{L},j\cdot R/2^{L})}|x_{k}-y|.$
Since the distance between each data point and $y$ is no more than the length of the interval $R/2^{L}$, and their total number is no more than $n$. When $L=\log(n)$, this error is $O(R)$.
Part II.
The second part is the Laplace noises. In our query, we add up $c$ and $s$ from $O(L)$ intervals. In these intervals, each contains a Laplace noise. Therefore, the total error is the sum of these noises:
$\mathbb{E}[|A-A^{\prime}|]\leq$ $\leavevmode\nobreak\ \mathbb{E}[|\sum_{i=1}^{L}\mathsf{Laplace}({RL}/{\epsilon})+y\cdot\sum_{i=1}^{L}\mathsf{Laplace}({L}/{\epsilon})|]$
$\leq$ $\leavevmode\nobreak\ \mathbb{E}[|\sum_{i=1}^{L}\mathsf{Laplace}({RL}/{\epsilon})|]+|y|\cdot\mathbb{E}[|\sum_{i=1}^{L}\mathsf{Laplace}({L}/{\epsilon})|]$
where the second step follows from triangle inequality.
For a random variable $Q$ with $\mathbb{E}[Q]=0$, using Fact 2.1 we have $\mathbb{E}[|Q|]\leq\sqrt{\mathop{\mathrm{Var}}[Q]}$, thus
$\mathbb{E}[|A-A^{\prime}|]\leq$ $\leavevmode\nobreak\ (\mathop{\mathrm{Var}}[\sum_{i=1}^{L}\mathsf{Laplace}({RL}/{\epsilon})])^{1/2}+|y|\cdot(\mathop{\mathrm{Var}}[\sum_{i=1}^{L}\mathsf{Laplace}({L}/{\epsilon})])^{1/2}$
$\leq$ $\leavevmode\nobreak\ \sqrt{L\cdot\mathop{\mathrm{Var}}[\mathsf{Laplace}({RL}/{\epsilon})]}+|y|\cdot\sqrt{L\cdot\mathop{\mathrm{Var}}[\mathsf{Laplace}({L}/{\epsilon})]}$
$=$ $\leavevmode\nobreak\ \sqrt{L\cdot{2R^{2}L^{2}}/{\epsilon^{2}}}+|y|\cdot\sqrt{L\cdot{2L^{2}}/{\epsilon^{2}}}$
$=$ $\leavevmode\nobreak\ \sqrt{2}\cdot{(R+|y|)L^{1.5}}/{\epsilon}$
$=$ $\leavevmode\nobreak\ O(\epsilon^{-1}R\log^{1.5}(n)).$
where the third step follows from $\mathop{\mathrm{Var}}[\mathsf{Laplace}(x)]=2x^{2}$.
The last step is because $y\in[0,R)$ and $L=\log(n)$. ∎
3.5 One Dimensional Differentially Private Data Structure
Collectively, we present the init time, space, query time, approximation error, and privacy guarantees for the 1-dimensional PD data structure (Algorithm 1, 2) in the next theorem. As a reminder, we address the 1-dimensional KDE query problem (Problem 1 with $d=1$) using the $\ell_{1}$ kernel distance.
Theorem 3.1 ($1$-Dimensional $\ell_{1}$ DP KDE Query).
Given a dataset $X\subset\mathbb{R}$ with $|X|=n$, there is an algorithm that uses $O(n)$ space to build a data structure (Algorithm 1, 2) which supports the following operations
  • Init$(X,\epsilon)$. It takes dataset $X$ and privacy parameter $\epsilon$ as input and spends $O(n)$ time to build a data-structure.
  • Query$(y\in\mathbb{R})$. It takes $y$ as input and spends $O(\log n)$ time to output a scalar $A$ such that $\mathbb{E}[|A-A^{\prime}|]\leq O(\epsilon^{-1}R\log^{1.5}n)$.
Furthermore, the data structure is $\epsilon$-DP.
Proof.
By Lemma 3.2, we prove the storage space, and the initialization time.
By Lemma 3.3, we prove the query time of the data-structure.
By Lemma 3.4, we prove the differential privacy guarantee.
By Lemma 3.5, we prove the error guarantee after adding the noise. ∎
3.6 High Dimensional $\ell_{1}$ Distance Query
We now generalize our 1-dimensional DP data structure to $d$-dimensional DP KDE queries. Specifically, we apply our 1-dimensional DP data structure independently to each dimension to compute high-dimensional queries.
Algorithm 3 High Dimensional Distance Query
1:Input: $n$ $d$-dimensional points in the box $[0,R)^{d}$, privacy parameter $\epsilon$, query $y$.
2:data structure HighDimFasterTree
3:members
4:     $T_{i}$ for $i\in[d]$triangleright $T_{i}$ represents a FasterTree in Algorithm 1 for $i-$th dimension.
5:end memebers
6:procedure Init($X$)
7:     Initialize $d$ independent data structure $T_{i}$ for $i\in[d]$ with Algorithm 1. $T_{i}$ builds on $i$-th coordinate projections of the data points. Each data structure is $\epsilon/d$-private.
8:end procedure
9:procedure Query($y$)
10:     Return the sum of the outputs of Algorithm 2 from $T_{i}$ with query $y_{i}$ for $i\in[d]$.
11:end procedure
12:end data structure
Lemma 3.6 (Differential Privacy).
The data structure HighDimFasterTree returned by HighDimFasterTree.Init (Algorithm 3) is $\epsilon$-DP.
Proof.
The proof directly follows from calling Lemma 3.4 for $d$ times with replacing $\epsilon$ by $\epsilon/d$. Then using the DP composition Lemma. ∎
Lemma 3.7 (Error Guarantee).
Let $X=\{x_{i}\}_{i\in[n]}$ be a dataset of $n$ $d$-dimensional vectors $x_{i}\in[0,R)^{d}$. Let $A^{\prime}=\sum_{i=1}^{n}|x_{i}-y|_{1}$ be the true distance query value. Let $A$ be the output of Query (Algorithm 3). Then we have $\mathbb{E}[|A-A^{\prime}|]\leq O(\epsilon^{-1}Rd^{1.5}\log^{1.5}n)$.
Proof.
The output of Algorithm 3 is the sum of each dimension. Let $A^{\prime}_{i}$ be the output of $D_{i}$, $A_{i}$ be the true distance of the $i$-th dimension. The total error of Algorithm 3 is $|\sum_{i=1}^{d}(A_{i}-A^{\prime}_{i})|$. From the proof of Algorithm 2, we know $A_{i}-A^{\prime}_{i}$ is the sum of $O(L)$ $\mathsf{Laplace}$ noises, where $L=O(\log n)$. Therefore, we have $\mathbb{E}[A_{i}-A^{\prime}_{i}]=0$, $\mathop{\mathrm{Var}}[A_{i}-A^{\prime}_{i}]=O(\epsilon^{-2}d^{2}R^{2}\log^{3}(n))$. Using the independence property of $A_{i}-A^{\prime}_{i}$ for all $i$, we bound $|\sum_{i=1}^{d}(A_{i}-A^{\prime}_{i})|$ by
$\mathbb{E}[|\sum_{i=1}^{d}(A_{i}-A^{\prime}_{i})|]\leq$ $\leavevmode\nobreak\ (\mathop{\mathrm{Var}}[\sum_{i=1}^{d}(A_{i}-A^{\prime}_{i})])^{1/2}$
$=$ $\leavevmode\nobreak\ (d\mathop{\mathrm{Var}}[A_{i}-A^{\prime}_{i}])^{1/2}$
$=$ $\leavevmode\nobreak\ O(d\cdot\epsilon^{-2}d^{2}R^{2}\log^{3}n)^{1/2}$
$=$ $\leavevmode\nobreak\ O(\epsilon^{-1}d^{1.5}R\log^{1.5}(n)).$
where the first step follows from Fact 2.1.
This completes the proof. ∎
Theorem 3.2 ($d$-Dimensional $\ell_{1}$ DP KDE Query, Formal version of Theorem 1.1).
Given a dataset $X\subset\mathbb{R}^{d}$ with $|X|=n$, there is an algorithm that uses $O(nd)$ space to build a data-structure (Algorithm 1, 2) which supports the following operations
  • Init$(X\subset\mathbb{R}^{d},\epsilon\in(0,1))$. It takes dataset $X$ and privacy parameter $\epsilon\in(0,1)$ as input and spends $O(nd)$ time to build a data structure.
  • Query$(y\in\mathbb{R}^{d})$. It takes $y$ as input and spends $O(d\log n)$ time to output a scalar $A\in\mathbb{R}$ such that $\mathbb{E}[|A-A^{\prime}|]\leq O(\epsilon^{-1}Rd^{1.5}\log^{1.5}n)$.
Proof.
By Lemma 3.6, we prove the privacy guarantees.
By Lemma 3.7, we prove the error guarantee of the Query. ∎
Acknowledgments
This work was completed during JH’s internship at Ensemble AI. JH would also like to thank Maojiang Su, Dino Feng and Andrew Chen for enlightening discussions on related topics.
JH is supported by Ensemble AI. HL is partially supported by NIH R01LM1372201. The content is solely the responsibility of the authors and does not necessarily represent the official views of the funding agencies.
Appendix
[…]
Appendix A Preliminaries
Fact A.1 (Union Bound).
Let $A_{1},A_{2},...,A_{n}$ be a set of probability events. It holds
$\Pr[\bigcup_{i=1}^{n}A_{i}]\leq\sum_{i=1}^{n}\Pr[A_{i}].$
Proof.
We prove union bound by induction. For $n=1$, $\Pr[A_{1}]=\Pr[A_{1}]$. Suppose this inequality holds true for $k$, then for $k+1$,
$\Pr[\bigcup_{i=1}^{k+1}A_{i}]=$ $\leavevmode\nobreak\ \Pr[\bigcup_{i=1}^{k}A_{i}]+\Pr[A_{k+1}]-\Pr[(\bigcup_{i=1}^{k}A_{i})\bigcap A_{k+1}]$
$\leq$ $\leavevmode\nobreak\ \Pr[\bigcup_{i=1}^{k}A_{i}]+\Pr[A_{k+1}]$
$\leq$ $\leavevmode\nobreak\ (\sum_{i=1}^{k}\Pr[A_{i}])+\Pr[A_{k+1}]$
$=$ $\leavevmode\nobreak\ \sum_{i=1}^{k+1}\Pr[A_{i}].$
Thus, we complete the proof. ∎
Appendix B $\ell_{2}$ Distance Query
Using the same technique in [BLM+24], we extend our $\ell_{1}$ result to $\ell_{2}$ norm. With our faster tree data structure, we reduce the error in [BLM+24] from $(1+\alpha,O(\epsilon^{-1}\alpha^{-1.5}R\log^{2}n))$ to $(1+\alpha,{O}(\epsilon^{-1}\alpha^{-1}R\log^{2}n))$. This method uses a mapping from $\ell_{2}$ to $\ell_{1}$ space that retains distances between data points with high probability.
References Approx. Ratio Error Query Time Init time Space
[BLM+24] $1+\alpha$ $\alpha^{-1.5}\epsilon^{-1}R\log^{2}n$ $\alpha^{-2}(d+\alpha^{-1}\log^{2}n)\log n\cdot\log(1/\alpha)$ $\alpha^{-2}nd\log n\cdot\log(1/\alpha)$ $n(d+\alpha^{-2}\log n\log(1/\alpha))$
Theorem B.1 $1+\alpha$ $\alpha^{-1}\epsilon^{-1}R\log^{2}n$ $\alpha^{-2}(d+\log n)\log n$ $\alpha^{-2}nd\log n$ $n(d+\alpha^{-2}\log n)$
Table 2: Comparison of $\norm{x-y}_{2}$ Results with Best Known Algorithm by [BLM+24]. We ignore the big-O notation in the table, for simplicity of presentation. We use $n$ to denote the number of points in dataset. We use $d$ to denote the dimension for each point in the dataset. We assume all the points are bounded by $R$. We use the $\epsilon$ to denote $\epsilon$-DP.
B.1 Error and DP Guarantees
Lemma B.1 ([Mat02]).
Let $\gamma\in(0,1)$ and define $T:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ by
$T(x)_{i}=\frac{1}{\beta k}\sum_{j=1}^{d}Z_{ij}x_{j},i\in[k],$
where $\beta=\sqrt{2/\pi}$ and $Z_{ij}$ are standard Gaussians. Then for every vector $x\in\mathbb{R}^{d}$, we have
$\Pr[(1-\gamma)\|x\|_{2}\leq\|T(x)\|_{1}\leq(1+\gamma)\|x\|_{2}]\geq 1-e^{-c\gamma^{2}k},$
where $c>0$ is a constant
Theorem B.1.
Let $X\subset\mathbb{R}^{d}$ be a private dataset of size $n$ with a bounded diameter of $R$ in $\ell_{2}$. For any $\alpha\in[0,1]$, there exists an $\epsilon$-DP data structure such that for any fixed query $y\in\mathbb{R}^{d}$, with probability $0.99$, it outputs $Z\in\mathbb{R}$ satisfying
$|Z-\sum_{x\in X}\|x-y\|_{2}|\leq\alpha\sum_{x\in X}\|x-y\|_{2}+{\widetilde{O}}(\epsilon^{-1}\alpha^{-1}R).$
Proof.
Let $\gamma=\alpha,$ and $k=O((\log n)/\alpha^{2})$. Our algorithm first maps $X$ to $T(X)$ using Lemma B.1, then solves $\ell_{1}$ query on $T(X)$ with Algorithm 3. By triangle inequality, we have
$\leavevmode\nobreak\ |Z-\sum_{x\in X}\|x-y\|_{2}|$
$\leq$ $\leavevmode\nobreak\ \underbrace{|\sum_{x\in X}\|T(x)-T(y)\|_{1}-\sum_{x\in X}\|x-y\|_{2}|}_{\mathrm{err}_{1}}+\underbrace{|Z-\sum_{x\in X}\|T(x)-T(y)\|_{1}|}_{\mathrm{err}_{2}}.$
Therefore, we divide the error into two parts, $\mathrm{err}_{1}$ and $\mathrm{err}_{2}$. $\mathrm{err}_{1}$ is from mapping $T$ (Lemma B.1), $\mathrm{err}_{2}$ is from Algorithm 3.
Bound of $\mathrm{err}_{1}$.
We prove with probability $0.99$,
$|\sum_{x\in X}\|T(x)-T(y)\|_{1}-\sum_{x\in X}\|x-y\|_{2}|\leq\alpha\sum_{x\in X}\|x-y\|_{2}.$
From Lemma B.1, let $\gamma=\alpha,k=c^{-1}(\log n+c^{\prime})/\alpha^{2}$, where $c^{\prime}$ is a sufficiently large constant. We have
$\Pr[|\|x\|_{2}-\|T(x)\|_{1}|>\alpha\|x\|_{2}]\leq e^{-(c^{\prime}+\log n)}\leq 0.01\cdot\frac{1}{n}.$
Therefore,
$\leavevmode\nobreak\ \Pr[|\sum_{x\in X}\|T(x)-T(y)\|_{1}-\sum_{x\in X}\|x-y\|_{2}|\leq\alpha\sum_{x\in X}\|x-y\|_{2}]$
$\geq$ $\leavevmode\nobreak\ \Pr[\;\forall x\in X\mid|\|x\|_{2}-\|T(x)\|_{1}|\leq\alpha\|x\|_{2}]$
$\geq$ $\leavevmode\nobreak\ 1-\sum_{x\in X}\Pr[\ |x\|_{2}-\|T(x)\|_{1}|>\alpha\|x\|_{2}]$
$\geq$ $\leavevmode\nobreak\ 1-n\cdot 0.01\cdot\frac{1}{n}$
$=$ $\leavevmode\nobreak\ 0.99$
The second step follows from union bound (Fact A.1).
Bound of $\mathrm{err}_{2}$.
After mapping $T$, with high probability, each coordinate ranges from 00 to $R^{\prime}:=O(R/k)$. Therefore, the additive error caused by Algorithm 3 equals
$|Z-\sum_{x\in X}\|T(x)-T(y)\|_{1}|=$ $\leavevmode\nobreak\ O(\epsilon^{-1}R^{\prime}k^{1.5}\log^{1.5}n)$
$=$ $\leavevmode\nobreak\ O(\epsilon^{-1}\frac{R\alpha^{2}}{\log n}(\frac{\log n}{\alpha^{2}})^{1.5}\log^{1.5}n)$
$=$ $\leavevmode\nobreak\ O(\alpha^{-1}\epsilon^{-1}R\log^{2}n),$
where the second step follows from choice of $R^{\prime}$.
This completes the proof. ∎
B.2 Init and Query Time
Finally, we prove the init time and query time of our algorithm for $\ell_{2}$ kernel distance.
Init Time.
For initializing, our algorithm contains two steps. First we map each data point from $d$-dimensional to $k$-dimensional. This step is a matrix multiplication and hence takes $O(ndk)$ time. Then we build a balanced binary tree. This step takes $O(nk)$ time. Therefore, given $k=(\log n)/\alpha^{2}$, the total init time is $O(\alpha^{-2}nd\log n)$.
Query Time.
For each query, our algorithm also contains two steps. First we map query $y$ from $d$-dimensional to $k$-dimensional. This step is a matrix multiplication and hence takes $O(dk)$ time. Then we query $y$ on the balanced binary tree. This step takes $O(k\log n)$ time. Therefore, given $k=(\log n)/\alpha^{2}$, the total query time is $O(\alpha^{-2}(d+\log n)\log n)$.
Appendix C $\ell_{p}^{p}$ Distance Query
Suppose $p\geq 1$ is an integer. To calculate $\ell_{p}^{p}$ distance queries, we use binomial expansion techniques. Similar to $\ell_{1}$, we first consider the one-dimensional case.
References Approx. Ratio Error Query Time Init time Space
[BLM+24] $1+\alpha$ $\epsilon^{-1}\alpha^{-0.5}R^{p}d^{1.5}\log^{1.5}n$ $\alpha^{-1}d\log^{2}n$ $nd$ $nd$
$\ell_{p}^{p}$ $1$ $\epsilon^{-1}p^{-0.5}R^{p}d^{1.5}\log^{1.5}n$ $d\log n$ $(n+pn^{1/p})d$ $pn^{1/p}d$
Table 3: Comparison of $\norm{x-y}_{p}^{p}$ Results with Best Known Algorithm by [BLM+24]. We ignore the big-O notation in the table, for simplicity of presentation. We use $n$ to denote the number of points in dataset. We use $d$ to denote the dimension for each point in the dataset. We assume all the points are bounded by $R$. We use the $\epsilon$ to denote $\epsilon$-DP.
Remark C.1.
We remark that for the small $p$ regime $p\in[1,2]$, our result (see Table 3) match the init time and space of [BLM+24], improve the approximation from $(1+\alpha)$ to 1, reduce the error by a factor of $\alpha^{-0.5}$, and reduce the query time by a factor of $\alpha^{-1}\log n$.
We first extend our key observation Lemma 3.1 with binomial expansion.
Lemma C.1.
For a collection of numbers $\{x_{1},x_{2},\cdots,x_{n}\}\subset\mathbb{R}$ and a number $y\in\mathbb{R}$. We define two sets
$S_{+}:=$$\leavevmode\nobreak\ \{k\in[n]\leavevmode\nobreak\ :\leavevmode\nobreak\ x_{k}>y\}$
$S_{-}:=$$\leavevmode\nobreak\ \{k\in[n]\leavevmode\nobreak\ :\leavevmode\nobreak\ x_{k}<y\},$
It holds
$\sum_{k=1}^{n}|x_{k}-y|^{p}=\sum_{j=0}^{p}{p\choose j}y^{p-j}((-1)^{p-j}\sum_{k\in S_{+}}x_{k}^{j}+(-1)^{j}\sum_{k\in S_{-}}x_{k}^{j}),$
where (${p\choose j}$ denotes the binomial coefficient that (${p\choose j}=\frac{p!}{j!(p-j)!}$.
Proof.
We show that
$\sum_{k=1}^{n}|x_{k}-y|^{p}=$ $\leavevmode\nobreak\ \sum_{x_{k}\in S_{+}}(x_{k}-y)^{p}+\sum_{x_{k}\in S_{-}}(y-x_{k})^{p}$
$=$ $\leavevmode\nobreak\ (\sum_{x_{k}\in S_{+}}\sum_{j=0}^{p}(-1)^{p-j}{p\choose j}x_{k}^{j}y^{p-j})+(\sum_{x_{k}\in S_{-}}\sum_{j=0}^{p}(-1)^{j}{p\choose j}x_{k}^{j}y^{p-j})$
$=$ $\leavevmode\nobreak\ \sum_{j=0}^{p}({p\choose j}(-1)^{p-j}y^{p-j}\sum_{k\in S_{+}}x_{k}^{j})+\sum_{j=0}^{p}({p\choose j}(-1)^{j}y^{p-j}\sum_{k\in S_{-}}x_{k}^{j})$
$=$ $\leavevmode\nobreak\ \sum_{j=0}^{p}{p\choose j}y^{p-j}((-1)^{p-j}\sum_{k\in S_{+}}x_{k}^{j}+(-1)^{j}\sum_{k\in S_{-}}x_{k}^{j}).$
Thus, we complete the proof. ∎
Lemma C.1 presents a decomposition of the final answer. To calculate the final answer, on each node, we store $\sum_{x_{k}\in I}x_{k}^{q}$ for all $q\in\mathbb{N},0\leq q\leq p$. $I$ denotes the interval that the node maintains. For each query, we define $s_{\text{left},q}$ and $s_{\text{right},q}$ as $\sum_{k\in S_{-}}x_{k}^{q}$ and $\sum_{k\in S_{+}}x_{k}^{q}$ respectively. Then we calculate them and construct the final answer.
Algorithm 4 $\ell_{p}^{p}$ Distance Query
1:data structure $\ell_{p}^{p}$ FasterTree $\triangleright$ Theorem C.1
2:members
3:     $T:=\{s_{l,j,q}\}$triangleright $s_{l,j,q}$ represents $\sum_{x_{k}\in[(j-1)\cdot\frac{R}{2^{l-1}},j\cdot\frac{R}{2^{l-1}})}x_{k}^{q}$ as defined above.
4:end members
5:
6:procedure Init($X,n,\epsilon,L$)
7:     for each $x_{k}$ in $X$ do
8:         find the leaf node $j$ of layer $L$ that $x_{k}$ is in the interval $I_{L,j}=[(j-1)\cdot\frac{R}{2^{L-1}},j\cdot\frac{R}{2^{L-1}})$
9:         for $q$ from 00 to $p$ do
10:              $s_{L,j,q}\leftarrow s_{L,j,q}+x_{k}^{q}$
11:         end for
12:     end for
13:     for each layer $l$ from $L-1$ to $1$ do
14:         for each node $j$ in layer $l$ do
15:              for $q$ from 00 to $p$ do
16:                  $s_{l,j,q}\leftarrow s_{l+1,2j-1,q}+s_{l+1,2j,q}$
17:              end for
18:         end for
19:     end for
20:     for $q$ from 00 to $p$ do
21:         Add independent noises drawn from $\mathsf{Laplace}({pLR^{q}}/{\epsilon})$ to each $s_{l,j,q}$
22:     end for
23:     $T\leftarrow\{s\}$
24:end procedure
25:
26:procedure Query($y\in\mathbb{R}$)
27:     for $q$ from 00 to $p$ do
28:         $s_{\text{left},q},s_{\text{right},q}\leftarrow 0$
29:     end for $\triangleright$ As previously defined
30:     for each layer $l$ from $2$ to $L$ do
31:         find the node $j$ of layer $l$ that $y$ is in the interval $I_{l,j}=[(j-1)\cdot\frac{R}{2^{l-1}},j\cdot\frac{R}{2^{l-1}})$
32:         for $q$ from 00 to $p$ do
33:              if $j\mod 2=0$ then $\triangleright$ if node $j$ is the right child of its parent
34:                  $s_{{\rm left},q}\leftarrow s_{{\rm left},q}+s_{l,j-1,q}$
35:              else $\triangleright$ if node $j$ is the left child of its parent
36:                  $s_{{\rm right},q}\leftarrow s_{{\rm right},q}+s_{l,j+1,q}$
37:              end if
38:         end for
39:     end for
40:     return $\sum_{q=0}^{p}{p\choose q}y^{p-q}((-1)^{p-q}s_{\text{right},q}+(-1)^{j}s_{\text{left},q})$triangleright Lemma C.1
41:end procedure
42:end data structure
Theorem C.1 ($1$-Dimensional $\ell_{p}^{p}$ DP KDE Query).
Given a dataset $X\subset\mathbb{R}$ with $|X|=n$, there is an algorithm that uses $O(pn^{1/p})$ space to build a data structure (Algorithm 4) which supports the following operations
  • Init$(X,\epsilon)$. It takes dataset $X$ and privacy parameter $\epsilon$ as input and spends $O(n+pn^{1/p})$ time to build a data-structure.
  • Query$(y\in\mathbb{R})$. It takes $y$ as input and spends $O(\log n)$ time to output a scalar $A$ such that $\mathbb{E}[|A-A^{\prime}|]\leq O(p^{-0.5}\epsilon^{-1}R^{p}\log^{1.5}n)$.
Furthermore, the data structure is $\epsilon$-DP.
Proof.
We set the total layers $L=(\log n)/p$.
Init Time.
The total number of nodes on the tree is $2^{L+1}-1=O(n^{1/p})$. Each node stores $p$ values, so there are $O(pn^{1/p})$ values stored on the tree. Plus the time of iterating all data points, initializing these values takes $O(n+pn^{1/p})$ time.
Query Time.
Each query iterate through all layers. On each layer it takes $O(p)$ time to calculate $s_{\text{left},q}$ and $s_{\text{right},q}$. There are $(\log n)/p$ layers, so the total query time is $O(\log n)$.
Privacy Guarantees.
Similar to Lemma 3.4, we consider each $q$ separately. For $q$, let $F_{q}(X):[0,R)^{n}\rightarrow\mathbb{R}^{2^{L+1}-1}$ be the mapping from dataset $X$ to $s_{l,j,q}$ of all $l,j$. Next we prove each $F_{q}(X)$ is $\epsilon/p$-DP.
For $F_{q}(X)$, when changing each data point, at most one node changes at each layer. Each node changes by at most $O(R^{q})$. Therefore, the sensitivity is $O(LR^{q})$. Adding coordinate-wise Laplace noise with magnitude $\eta=O(pLR^{q}/\epsilon)$ suffices to ensure $(\epsilon/p)$-DP using the standard Laplace mechanism.
By Lemma 2.1, the differential privacy parameter $\epsilon$ of the tree $T:=(F_{q}(X):0\leq q\leq p)$ equals $p\cdot\epsilon/p=\epsilon$. This completes the proof.
Error Guarantees.
Similar to Lemma 3.5, the additive error consists of two parts.
The first part is from the data in the leaf node which contains query $y$. The error is
$\sum_{x_{k}\in[(j-1)\cdot R/2^{L},j\cdot R/2^{L})}|x_{k}-y|^{p}\leq n\cdot(\frac{R}{2^{L}})^{p}.$
When $L=(\log n)/p$, this error is $O(R^{p})$.
The second part is the Laplace noise. We first show that
$\mathbb{E}[|\sum_{i=1}^{L}\mathsf{Laplace}(\lambda)|]\leq$ $\leavevmode\nobreak\ \sqrt{\mathop{\mathrm{Var}}[\sum_{i=1}^{L}\mathsf{Laplace}(\lambda)]}$
$\leq$ $\leavevmode\nobreak\ \sqrt{L\cdot\mathop{\mathrm{Var}}[\mathsf{Laplace}(\lambda)]}$
$=$ $\leavevmode\nobreak\ \sqrt{2L\lambda^{2}}$
$=$ $\leavevmode\nobreak\ \sqrt{2L}\lambda,$
where the first step follows from Fact 2.1, the third step follows from $\mathop{\mathrm{Var}}[\mathsf{Laplace}(x)]=2x^{2}$.
Replacing $\lambda=pR^{p}L/\epsilon$, we have
$\mathbb{E}[|\sum_{i=1}^{L}\mathsf{Laplace}(pR^{p}L/\epsilon)|]\leq\sqrt{2}pR^{p}L^{1.5}/\epsilon.$ (C.1)
Then we bound the error with this inequality:
$\mathbb{E}[|A-A^{\prime}|]\leq$ $\leavevmode\nobreak\ \mathbb{E}[|\sum_{q=0}^{p}{p\choose q}y^{p-q}\sum_{i=1}^{L}((-1)^{p-j}\mathsf{Laplace}({pR^{q}L}/{\epsilon})+(-1)^{j}\mathsf{Laplace}({pR^{q}L}/{\epsilon}))|]$
$\leq$ $\leavevmode\nobreak\ \sum_{q=0}^{p}{p\choose q}y^{p-q}\mathbb{E}[|\sum_{i=1}^{L}(\mathsf{Laplace}({pR^{q}L}/{\epsilon})+\mathsf{Laplace}({pR^{q}L}/{\epsilon}))|]$
$=$ $\leavevmode\nobreak\ \sum_{q=0}^{p}{p\choose q}y^{p-q}\cdot 2pR^{q}L^{1.5}$
$=$ $\leavevmode\nobreak\ 2pL^{1.5}\sum_{q=0}^{p}{p\choose q}y^{p-q}R^{q}$
$=$ $\leavevmode\nobreak\ 2pL^{1.5}(y+R)^{p}$
$=$ $\leavevmode\nobreak\ O(p^{-0.5}R^{p}\log^{1.5}n),$
where the third step follows from Eq. (C.1). The last step is from $L=(\log n)/p,y\in[0,R)$. ∎
For high dimensional $\ell_{p}^{p}$ queries, we use the same procedures in Section 3.6. We build $d$ independent data structures for each dimension following Algorithm 3. We also compare the privacy and error guarantees with those in [BLM+24] in Table 3.
Theorem C.2 ($d$-Dimensional $\ell_{p}^{p}$ DP KDE Query).
Given a dataset $X\subset\mathbb{R}^{d}$ with $|X|=n$, there is an algorithm that uses $O(pn^{1/p}d)$ space to build a data-structure which supports the following operations
  • Init$(X,\epsilon)$. It takes dataset $X$ and privacy parameter $\epsilon$ as input and spends $O(nd+pn^{1/p}d)$ time to build a data-structure.
  • Query$(y\in\mathbb{R}^{d})$. It takes $y$ as input and spends $O(d\log n)$ time to output a scalar $A$ such that $\mathbb{E}[|A-A^{\prime}|]\leq O(p^{-0.5}\epsilon^{-1}R^{p}d^{1.5}\log^{1.5}n)$.
Furthermore, the data structure is $\epsilon$-DP.
Proof.
We omit the proof as it follows from Section 3.6 in a straightforward manner. ∎